4 Question 20 A baseball is hit upward and travels along a parabolic arc before it strikes the ground. Which one of the following statements is necessarily true? The x-component of the velocity of the ball varies throughout the balls flight The acceleration of the ball decreases as the ball moves upward The y-component of the velocity of the ball is zero when the ball is at the highest point The velocity of the ball is a maximum when the ball is at the highest point in the arc The acceleration of the ball is zero when the ball is at the highest point in the arc - Question 2:1 4 pts spacecraft is traveling with a velocity of +4349.7 m/s. Suddenly the retrorockets are fred, and the ft begins to slow down with an acceleration whose magnitude is -11.235 m/s2 How far raft travel until it comes to a momentary halt B.42 x10 m 6.13 x10s m 9.46 x 10 m 3.17 x 10 m 5.28x 105 m
Answer #1

21) using third equation of motion,

s = v^2/2a

= 4349.7^2 /(2*11.235)

= 8.42*10^5 m

First option is correct